Step by step
The question
Question: x can be 0, 1, 2, 3, with probabilities 0.2, 0.3, a missing one, 0.1: find the missing probability.
x: 0, 1, 2, 3; P(x): 0.2, 0.3, ?, 0.1
1 All P(x) add to 1
x always takes one of its values.
So all the probabilities add up to 1.
a probability table always adds up to 1
0.2 plus 0.3 plus P of 2 plus 0.1 is 1.
0.2 + 0.3 + P(2) + 0.1 = 1
The known ones add up to 0.6.
0.6 + P(2) = 1
To remove the plus 0.6, we take away 0.6 on both sides.
1 minus 0.6 is 0.4.
P(2) = 1 − 0.6 = 0.4
Careful!
The missing probability is what is left over from 1.
Careful: the missing P is what is left of 1
2 Check
Let's check: 0.2 plus 0.3 plus 0.4 plus 0.1 is 1.
So it is right!
0.2 + 0.3 + 0.4 + 0.1 = 1
Let's look back.
First: all P(x) add to 1.
Last: check.
Answer
Careful: the missing P is what is left of 1